OpenStudy (anonymous):

Choose the correct simplification of the expression (xyz2)4. x5y5z6 x5y5z8 xyz16 x4y4z8 Yes they are to the the powers.

OpenStudy (anonymous):

4

OpenStudy (anonymous):

\[(xyz^2)^4 = (x)^4(y)^4(z^2)^4 = x^4y^4z^{2*4} = x^4y^4z^8\]